# Strict comparison in reduced group $C^{*}$-algebras

### v2

@@ 1.5. Comments
The inroad to prove strict comparison is to prove the more general free independence theorem in the norm ultrapower of the reduced group $C^{*}$ algebra, which is the selflessness property from [56]. Our inspiration for the argument to achieve this came from a recent very beautiful paper by Louder and Magee [44]. Curiously, our original approach to solving the problem of strict comparison was (motivated in a sense by the work of the third author [41]) to use strong convergence methods in random matrix theory, and this is what indirectly led us to [44]. Now, in order to prove this free independence theorem, one benefits by using the rapid decay property (see [58, 33]), which allows one to control the operator norm of elements in the group ring in terms of the radius of the ball of support and the 2-norm. A key next step is a quantitative analysis of the eventual faithfulness of certain tailored retraction maps that one constructs using the group structure; for instance, free product structure, or, more generally, an acylindrical action on a hyperbolic space. This is used to obtain a particularly strong existential inclusion of the group into the free product of itself with a copy of $\mathbb{Z}$. Piecing these ideas together allows one to push this inclusion to the level of the reduced group $C^{*}$-algebra, as needed.
A note to the reader
@@
The proofs of these results split into a hands-on computational part (showing selflessness in our sense, see 3.1) and a soft $C^{*}$-algebra part (using rapid decay à la Louder-Magee, see Theorem 3.5). If the reader would like to just focus on understanding the proof of Theorem A, we direct them to a self contained approach via the combination of Proposition 3.2 and Theorem 3.5. The quantitative part of the proof of Theorem B is rather involved, but quite natural from a geometric point of view. Appendix A contains a quantitative version of a key theorem of Wenyuan Wang that is needed to prove our result B. This is quite technical and we recommend the interested reader to side by side refer to [68, Corollary 3.4].
@@ Acknowledgments
This result was proved during the visit of the first author to UCSD in Fall 2024, supported by the third author’s NSF grant DMS 2350049 and the Institute of Mathematics of the Polish Academy of Sciences. The second and fourth author were supported in part by NSF grant DMS 2153805. We thank our colleagues J. Gabe, D. Jekel, N.C. Phillips, L. Robert, M. Rørdam, C. Schafhauser, G. Szabo, H. Thiel and S. White for extremely helpful comments on the introduction. We sincerely thank all the experts we shared our earlier draft with, for providing us great encouragement and helpful comments to improve our paper. We also thank our colleagues J. Behrstock, F. Fournier-Facio, J. Huang, Y. Lodha, D. Osin, A. Sisto, and W. Yang for helpful pointers.
@@ 2. Notation
We call these norms on ${\mathbb{C}}\Gamma$ the $\ell^{1}$-norm, the $\ell^{2}$-norm, and the operator norm respectively. As suggested by the definition of the operator norm, each element of ${\mathbb{C}}\Gamma$ can be viewed as a bounded operator on $\ell^{2}(\Gamma)$. We define the reduced group $C^{*}$-algebra $C_{r}^{*}(\Gamma)$ to be the operator norm closure of ${\mathbb{C}}\Gamma$ inside of ${\mathbb{B}}(\ell^{2}(\Gamma))$.
A state on a unital $C^{*}$-algebra $A$ is a positive linear functional $\phi$ such that $\phi(1)=1.$ We say a state $\phi$ is faithful if $\phi(x^{*}x)>0$ for any nonzero $x\in A$. We call $\phi$ a trace if $\phi(xy)=\phi(yx)$ for all $x,y\in A$. Let $A_{1},A_{2}\subset(A,\phi)$ be subalgebras. We say that $A_{1}$ and $A_{2}$ are freely independent relative to $\phi$ if whenever $a_{k}\in A_{i_{k}}$ with $\phi(a_{k})=0$ and $i_{k}\neq i_{k+1}$ for $1\leq k\leq n$ then $\phi(a_{1}\cdots a_{n})=0.$ If $(B_{1},\rho),(B_{2},\rho)$ are unital $C^{*}$-algebras with faithful states, then there is a unital $C^{*}$-algebra with faithful state $(B,\rho)$ and embeddings $\pi_{i}:(B_{i},\rho)\hookrightarrow(B,\rho)$ such that $\pi_{1}(B_{1})$ and $\pi_{2}(B_{2})$ are freely independent relative to $\rho.$ We call $(B,\rho)$ the reduced free product of $(B_{1},\rho_{1})$ and $(B_{2},\rho_{2})$; we omit the states when they are clear from context (see [4, 63]).
@@
where $N_{\mathcal{U}}$ consists of all sequences which are eventually constant with respect to ${\mathcal{U}}$. More precisely, a sequence $(g_{n})_{n}\in N$ if and only if there is $h\in\Gamma$ and $S\in{\mathcal{U}}$ such that $g_{n}=h$ for all $n\in S.$ We call $\Gamma^{\mathcal{U}}$ an ultrapower of $\Gamma$. If $A$ is a $C^{*}$-algebra, we define $A^{\mathcal{U}}$ as
@@
where $I_{\mathcal{U}}$ consists of all norm-bounded sequences $(x_{n})$ in $A$ with the following property: for all $\varepsilon>0$ there is $S\in{\mathcal{U}}$ such that for all $n\geq N$, $\|x_{n}\|<\varepsilon$. We call $A^{\mathcal{U}}$ an ultrapower of $A$.
@@ 3.1. A new property for groups
The following is a quantitative strengthening of the definition of the mixed identity free (MIF) property for groups [37], and our terminology is motivated by [56].
@@
We say a group with a finite generating set $(G,X)$ is selfless if there is a function $f:{\mathbb{N}}\to{\mathbb{R}}$ with $\liminf_{n}f(n)^{\frac{1}{n}}=1$ such that for all $n\geq 1$, there is an epimorphism $\phi_{n}:G*\langle a\rangle\to G$ such that $\phi_{n}|_{G}=\operatorname{id}_{G}$, $\phi_{n}$ is injective on $B_{X\cup\{a\}}(n)$, and $\phi_{n}(B_{X\cup\{a\}}(n))\subset B_{X}(f(n))$. Here, $a$ is a torsion-free generator so that $G*\langle a\rangle\cong G*{\mathbb{Z}}$.
@@
If $(G,X)$ contains a torsion-free element and $(H,Y)$ is an infinite group, then $G*H$ is selfless. In particular, ${\mathbb{F}}_{n}$ is selfless for all $n\geq 2.$
@@
Let $g\in G$ be torsion-free. Set $p=|g|_{X}$. For each $n$ let $h_{n}\in H$ be an element such that $|h_{n}|_{Y}>2n$ and $\liminf_{n}|h_{n}|_{Y}^{\frac{1}{n}}=1.$ Define $\phi_{n}:G*H*\langle a\rangle\to G*H$ by $\phi_{n}|_{G*H}=\operatorname{id}_{G*H}$ and $\phi_{n}(a)=h_{n}gh_{n}^{-1}$. It is clear that $\liminf_{n}|\phi_{n}(a)|_{X\cup Y}^{\frac{1}{n}}=1$, and therefore
@@
for some function $f$ where $\liminf_{n}f(n)^{\frac{1}{n}}=1.$
@@
Now, assume $m>2$. Write all $s_{i}$ in standard form for the free product $G\ast H$. Then all the words in $s_{i}$ from $H$ have word length (with respect to $Y$) at most $2n$, while $|h_{n}|_{Y}>2n$, so no full cancellation can happen between them. Thus, if all $s_{i}$ with $1<i<m$ have terms in $G$, then the standard form of $s_{1}t_{1}\cdots s_{m-1}t_{m-1}s_{m}$ is,
$h^{\prime}_{1}g^{p_{1}}h^{\prime}_{2,1}g^{\prime}_{2,1}h^{\prime}_{2,2}\cdots g^{\prime}_{2,k_{2}-1}h^{\prime}_{2,k_{2}}g^{p_{2}}\cdots h^{\prime}_{m-1,k_{m-1}}g^{p_{m-1}}h^{\prime}_{m}$
where all $h^{\prime}$ are non-identity elements of $H$ and all $g^{\prime}$ are non-identity elements of $G$. In particular, it is not the identity. Therefore, there is some $1<i<m$ s.t. $s_{i}\in H$. But as $s_{i}\neq e$ for $1<i<m$, we have $h_{n}^{-1}s_{i}h_{n}$ is a non-identity element of $H$ when $s_{i}\in H$, so examining the standard form again, we see that such terms $h_{n}^{-1}s_{i}h_{n}$ will appear in the standard form, again a contradiction. This shows $m\leq 2$. Now, if $m=2$, then $s_{1}t_{1}\cdots s_{m-1}t_{m-1}s_{m}=s_{1}t_{1}s_{2}$. Again, any $H$-term in the standard forms of either $s_{1}$ or $s_{2}$ cannot fully cancel with $h_{n}$ or $h_{n}^{-1}$. So, $g^{p_{1}}$ will appear in the standard form, which is a contradiction. Hence, $m=1$ and $s_{1}=e$. ∎
@@ 3.2. A general family of selfless groups
We will prove here that all acylindrically hyperbolic groups with trivial finite radical are selfless. For this subsection, for background notation and preliminaries on acylindrically hyperbolic groups, please refer to the beautiful memoir [20], and also [48]. At the outset we wish to express our gratitude to the following two experts for their guidance. Firstly we sincerely thank D. Osin who first pointed out to us that our selflessness for free products (Proposition 3.2) will hold in the general setting of acylindrically hyperbolic groups. Secondly we are very grateful to W. Yang for kindly sharing with us his ideas towards a different approach. First we describe the outline for our proof:
@@
Let $G$ be an acylindrically hyperbolic group with trivial finite radical. Then $G$ is selfless.
@@
Fix a finite generating set $S$ for $G$. Since $G$ is acylindrically hyperbolic and has trivial finite radical, by [20, Theorem 6.14], there is a hyperbolically embedded copy of ${\mathbb{F}}_{2}$ in $G$. Denote by $a,b$ the generators of ${\mathbb{F}}_{2}$. Since $\langle a\rangle$ is hyperbolically embedded in ${\mathbb{F}}_{2}$, by transitivity of hyperbolic embeddings [20, Proposition 4.35] we have that $\langle a\rangle$ is hyperbolically embedded in $G$. Therefore $\langle a\rangle$ is almost malnormal in $G$ [20, Proposition 4.33], implying that the maximal virtually cyclic subgroup containing $a,$ namely $E(a)$, is equal to $\langle a\rangle$ (see [26] for details on virtually cyclic groups). By [20, Theorem 4.31], we have that this copy of ${\mathbb{F}}_{2}$, with word metric given by its generating set $\{a,b\}$, is a Lipschitz quasi-retract of $G$ with word metric given by $S$. Therefore there are constants $C,D$ such that whenever $g\in{\mathbb{F}}_{2},$ $|g|_{\{a,b\}}\leq C|g|_{S}$ and $|g|_{S}\leq D|g|_{\{a,b\}}$. Since $a$ has infinite order, [48, Theorem 1.4] says that there is an acylindrical action of $G$ on a hyperbolic space $X$ such that $a$ is loxodromic.
Now consider the sequence $g_{n}=b^{n}ab^{-n}$. We note that each $g_{n}$ is loxodromic for the action $G\curvearrowright X$; moreover, each $g_{n}$ is primitive. Thus $E(g_{n})=\langle g_{n}\rangle$ for all $n.$ We note that for all $n\geq\frac{1}{2}C^{2}D$ and any $h\in B_{S}({|g_{n}|_{S}^{1/2}})$ (the $|g_{n}|_{S}^{1/2}$-ball around the identity in the Cayley graph of $G$ with the generating set $S$), we have $h\not\in E(g_{n})$. Indeed, $|g_{n}|_{\{a,b\}}=2n+1>C^{2}D$ which implies that either $h\not\in{\mathbb{F}}_{2}$ or
@@
while clearly every element $g^{\prime}\in E(g_{n})$ is an element of ${\mathbb{F}}_{2}$ with $|g^{\prime}|_{\{a,b\}}\geq|g_{n}|_{\{a,b\}}$.
Since all of the $g_{n}$ are conjugate, they all have equal stable length. Therefore we can apply Theorem A.4 to get a degree 6 polynomial $P$ such that whenever $h_{1},\ldots,h_{m}\in B_{S}({|g_{n}|_{S}^{1/2}})$ and $|k_{1}|,\ldots,|k_{m}|\geq P(|g_{n}|_{S})$ we have $h_{1}g_{n}^{k_{1}}\cdots h_{m}g_{n}^{k_{m}}\neq e$. In particular, taking $g_{n}^{\prime}=g_{n}^{P(|g_{n}|_{S})}$ we have $h_{1}(g_{n}^{\prime})^{k_{1}}\cdots h_{m}(g_{n}^{\prime})^{k_{m}}\neq e$ for any non-zero integers $k_{i}$. This, along with ${\mathbb{F}}_{2}$ being a Lipschitz quasi-retract of $G$, implies that there is a sequence $g_{n}^{\prime\prime}\in G$ with $|g_{n}^{\prime\prime}|_{S}=O(n^{12})$ and a constant $N$ so that for any $n>N$, we have that any word $w=h_{1}(g_{n}^{\prime\prime})^{n_{1}}h_{2}(g_{n}^{\prime\prime})^{n_{2}}\cdots h_{k}(g_{n}^{\prime\prime})^{n_{k}}h_{k+1}\neq e$ where $h_{i}\in B_{S}(n)$, concluding the proof.

### v4

@@ 1.4. Comments to the reader
The inroad to prove strict comparison is to prove the more general free independence theorem in the norm ultrapower of the reduced group $C^{*}$-algebra, which is the selflessness property from [59]. As mentioned previously, our inspiration for the argument to achieve this came from a recent very beautiful paper by Louder and Magee [46]. Curiously, our original approach to solving the problem of strict comparison was (motivated in a sense by the work of the third author [43]) to use strong convergence methods in random matrix theory, and this is what indirectly led us to [46]. Now, in order to prove this free independence theorem, one benefits by using the rapid decay property (see [61, 35]), which allows one to control the operator norm of elements in the group ring in terms of the radius of the ball of support and the 2-norm. A key next step is a quantitative analysis of the eventual faithfulness of certain tailored retraction maps that one constructs using the group structure; for instance, free product structure, or, more generally, an acylindrical action on a hyperbolic space. This is used to obtain a particularly strong existential inclusion of the group into the free product of itself with a copy of $\mathbb{Z}$. Piecing these ideas together allows one to push this inclusion to the level of the reduced group $C^{*}$-algebra, as needed.
Concerning the organization
@@
The proofs of these results split into a hands-on computational part (showing selflessness in our sense, see Definition 3.1) and a soft $C^{*}$-algebra part (using rapid decay à la Louder-Magee, see Theorem 3.5), and then reaching the finish line for strict comparison with Robert’s [59]. If the reader would like to just focus on understanding the proof for arbitrary free products with rapid decay, we direct them to a self contained approach via the combination of Proposition 3.2 and Theorem 3.5. We also remark to the reader that after our paper was released publicly, M. Magee kindly shared with us his proof (via a quantitative Baumslag lemma) for just the non abelian free group case. This is in Subsection 3.4.
On the other hand, the proof of Theorem B is quite involved, but the quantitative part is somewhat natural from a geometric point of view. Appendix A contains a quantitative version of a key theorem of Wenyuan Yang that is needed to prove our result B. This is quite technical and we recommend the interested reader to side by side refer to [72, Corollary 3.4].
@@ Acknowledgments
This result was proved during the visit of the first author to UCSD in Fall 2024, supported by the third author’s NSF grant DMS 2350049 and the Institute of Mathematics of the Polish Academy of Sciences. The second and fourth author were supported in part by NSF grant DMS 2153805. We thank our colleagues J. Gabe, D. Jekel, N.C. Phillips, L. Robert, M. Rørdam, C. Schafhauser, G. Szabo, H. Thiel and S. White for extremely helpful comments on the introduction. We sincerely thank all the experts we shared our earlier draft with, for providing us great encouragement and helpful comments to improve our paper. We also thank our colleagues J. Behrstock, F. Fournier-Facio, J. Huang, Y. Lodha, M. Magee, D. Osin, A. Sisto, and W. Yang for helpful pointers and feedback. We thank the anonymous referee for several very helpful and detailed comments that improved the paper.
@@ 2. Notation
We call these norms on ${\mathbb{C}}\Gamma$ the $\ell^{1}$-norm, the $\ell^{2}$-norm, and the reduced operator norm respectively. As suggested by the definition of the operator norm, each element of ${\mathbb{C}}\Gamma$ can be viewed as a bounded operator on the Hilbert space $\ell^{2}(\Gamma)$, by the left regular representation $\lambda_{\Gamma}:{\mathbb{C}}\Gamma\to\mathbb{B}(\ell^{2}\Gamma)$ (define $\lambda_{\Gamma}(g)(\delta_{h})=\delta_{gh}$ and extend linearly). Note therefore we can write $\|x\|=\|\lambda_{\Gamma}(x)\|$ when viewed as a bounded operator on $\ell^{2}(\Gamma)$. Note that for any element $x\in\mathbb{C}\Gamma$, one has the following elementary inequality $\|x\|_{2}\leq\|x||\leq\|x\|_{1}.$ We define the reduced group $C^{*}$-algebra $C_{r}^{*}(\Gamma)$ to be the operator norm closure of ${\mathbb{C}}\Gamma$ inside of ${\mathbb{B}}(\ell^{2}(\Gamma))$. We will also need the following elementary fact, which follows from the triangle inequality: if $\phi:\Gamma\to G$ is a group homomorphism then $\phi$ extends to an $\ell^{1}$-contractive map between ${\mathbb{C}}\Gamma$ and ${\mathbb{C}}G$.
A state on a unital $C^{*}$-algebra $A$ is a positive linear functional $\phi$ such that $\phi(1)=1.$ We say that a state $\phi$ is faithful if $\phi(x^{*}x)>0$ for any nonzero $x\in A$. We call $\phi$ a trace if $\phi(xy)=\phi(yx)$ for all $x,y\in A$. Let $A_{1},A_{2}\subset(A,\phi)$ be subalgebras. We say that $A_{1}$ and $A_{2}$ are freely independent relative to $\phi$ if whenever $a_{k}\in A_{i_{k}}$ with $\phi(a_{k})=0$ and $i_{k}\neq i_{k+1}$ for $1\leq k<n$ then $\phi(a_{1}\cdots a_{n})=0.$ If $(B_{1},\rho_{1}),(B_{2},\rho_{2})$ are unital $C^{*}$-algebras with faithful states, then there is a unital $C^{*}$-algebra with a faithful state $(B,\rho)$ and embeddings $\pi_{i}:(B_{i},\rho_{i})\hookrightarrow(B,\rho)$ such that $\pi_{1}(B_{1})$ and $\pi_{2}(B_{2})$ are freely independent relative to $\rho.$ We call $(B,\rho)$ the reduced free product of $(B_{1},\rho_{1})$ and $(B_{2},\rho_{2})$; we omit the states when they are clear from context (see [5, 66]), and write $B_{1}*B_{2}$. We have in this context that whenever $G_{1}$ and $G_{2}$ are countable groups, $C^{*}_{r}(G_{1}*G_{2})\cong C^{*}_{r}(G_{1})*C^{*}_{r}(G_{2})$, see Proposition 1.5.3 in [67].
@@
where $N_{\mathcal{U}}$ consists of all sequences which are eventually identity with respect to $\mathcal{U}$}. More precisely, a sequence $(g_{n})_{n}\in N_{\mathcal{U}}$ if and only if there is $S\in{\mathcal{U}}$ such that $g_{n}=1_{\Gamma}$ for all $n\in S.$ We call $\Gamma^{\mathcal{U}}$ an ultrapower of $\Gamma$. If $A$ is a $C^{*}$-algebra, we define $A^{\mathcal{U}}$ as
@@
where $I_{\mathcal{U}}$ is the ideal consisting of all norm-bounded sequences $(x_{n})$ in $A$ with the following property: for all $\varepsilon>0$ there is $S\in{\mathcal{U}}$ such that for all $n\in S$, $\|x_{n}\|<\varepsilon$. We call $A^{\mathcal{U}}$ an ultrapower of $A$. Note that under the definition $\|(x_{n})_{\mathcal{U}}\|_{A^{\mathcal{U}}}=\lim_{\mathcal{U}}\|x_{n}\|$, and the natural product structure, $A^{\mathcal{U}}$ is a $C^{*}$-algebra.
@@ 3.1. A new property for groups
The following is a quantitative strengthening of the definition of the mixed identity free (MIF) property for groups [40], and our terminology is motivated by [59] and is very much related to the quantitative Baumslag lemma from [46].
@@
We say a group with a finite generating set $(G,X)$ is selfless if there is a function $f:{\mathbb{N}}\to{\mathbb{R}}_{+}$ with $\liminf_{n}f(n)^{\frac{1}{n}}=1$ such that for all $n\geq 1$, there is an epimorphism $\phi_{n}:G*\langle a\rangle\to G$ such that $\phi_{n}|_{G}=\operatorname{id}_{G}$, $\phi_{n}$ is injective on $B_{X\cup\{a\}}(n)$, and $\phi_{n}(B_{X\cup\{a\}}(n))\subset B_{X}(f(n))$. Here, $a$ is a torsion-free generator so that $G*\langle a\rangle\cong G*{\mathbb{Z}}$.
@@
If $(G,X)$ contains a torsion-free element and $(H,Y)$ is an infinite group, then GHG*H is selfless. In particular, ${\mathbb{F}}_{n}$ is selfless for all $n\geq 2.$
@@
Let $g\in G$ be torsion-free. Set $p=|g|_{X}$. For each $n$ let $h_{n}\in H$ be an element such that $|h_{n}|_{Y}=2n+1$. Define $\phi_{n}:G*H*\langle a\rangle\to G*H$ by $\phi_{n}|_{G*H}=\operatorname{id}_{G*H}$ and $\phi_{n}(a)=h_{n}gh_{n}^{-1}$. It is clear that $\lim_{n}|\phi_{n}(a)|_{X\cup Y}^{\frac{1}{n}}=1$, and therefore
@@
for some function $f$ where $\lim_{n}f(n)^{\frac{1}{n}}=1.$
@@
Assume that $m>1$. We expand out the expression $s_{1}t_{1}\cdots s_{m-1}t_{m-1}s_{m}$ and regroup the product as follows:
$(s_{1}h_{n})g^{p_{1}}(h_{n}^{-1}s_{2}h_{n})g^{p_{2}}(h_{n}^{-1}s_{3}\cdots s_{m-1}h_{n})g^{p_{m-1}}(h_{n}^{-1}s_{m}).$
We claim that each term in parentheses begins and ends with a letter in $H$. Indeed, for each $s_{i}\in H$ (necessarily nontrivial) this is immediate. Otherwise, $s_{i}$ begins either with a letter in $G$ or a letter in $H$ of length at most $2n$ followed by a letter in $G$; $s_{i}$ ends in a similar way. Therefore $h_{n}^{-1}s_{i}h_{n}$ begins and ends with a letter in $H$. Since each $p_{i}$ is nonzero, we now see that the above product is a nontrivial word in the free product; this finishes the proof. ∎
@@ 3.2. A general family of selfless groups
We will prove here that all finitely generated acylindrically hyperbolic groups with trivial finite radical are selfless. For this subsection, for background notation and preliminaries on acylindrically hyperbolic groups, please refer to the beautiful memoir [22], and also [50]. At the outset we wish to express our gratitude to the following two experts for their guidance. Firstly we sincerely thank D. Osin who first pointed out to us that our selflessness for free products (Proposition 3.2) will hold in the general setting of acylindrically hyperbolic groups. Secondly we are very grateful to W. Yang for kindly sharing with us his ideas towards a different approach. First we describe the outline for our proof:
@@
Let $G$ be a finitely generated acylindrically hyperbolic group with trivial finite radical. Then $G$ is selfless.
@@
Fix a finite generating set $S$ for $G$. Since $G$ is acylindrically hyperbolic and has trivial finite radical, by [22, Theorem 6.14], there is a hyperbolically embedded copy of ${\mathbb{F}}_{2}$ in $G$. Denote by $a,b$ the generators of ${\mathbb{F}}_{2}$. Without loss of generality, assume $a,b\in S$. By the proof of [22, Theorem 4.31], there is a constant $K>0$ and a finite generating set $Y$ of ${\mathbb{F}}_{2}$ such that for all $h\in{\mathbb{F}}_{2}$, $|h|_{Y}\leq|h|_{S}$ and $|h|_{S}\leq K|h|_{Y}$. Since $\{a,b\}$ and $Y$ are both finite generating sets for ${\mathbb{F}}_{2}$, we have that $|h|_{a,b}$ and $|h|_{Y}$ are quasi-equivalent. Therefore there are constants $C,D$ such that whenever $h\in{\mathbb{F}}_{2},$ $|h|_{\{a,b\}}\leq C|h|_{S}$ and $|h|_{S}\leq D|h|_{\{a,b\}}$.
Set $g_{n}=b^{n}ab^{-n}$. We note that each $g_{n}$ is primitive. Note that for all $n,$ ${\mathbb{F}}_{2}=\langle b\rangle*\langle g_{n}\rangle$. By hyperbolicity of free factors ([22, Example 2.2(c)]) and transitivity of hyperbolic embeddings ([22, Proposition 4.35]) we have that $\langle g_{n}\rangle$ is hyperbolically embedded in $G$. In particular, since $a$ has infinite order and is contained inside a hyperbolically embedded cyclic subgroup, [50, Theorem 1.4] says that there is an acylindrical action of $G$ on a hyperbolic space $X$ such that $a$ is loxodromic. It follows that $g_{n}$ is loxodromic for each $n$ too. By almost malnormality of hyperbolically embedded subgroups ([22, Proposition 4.33]), [22, Remark 6.2], and [22, Corollary 6.6], it follows that $E(g_{n})=\langle g_{n}\rangle$ for each $n$.
We note that for all $n\geq\frac{1}{2}C^{2}D$ and any $h\in B_{S}({|g_{n}|_{S}^{1/2}})\setminus\{e\}$ (the $|g_{n}|_{S}^{1/2}$-ball around the identity in the Cayley graph of $G$ with the generating set $S$), we have $h\not\in E(g_{n})$. Indeed, $|g_{n}|_{\{a,b\}}=2n+1>C^{2}D$ which implies that either $h\not\in{\mathbb{F}}_{2}$ or
@@
while clearly every element $g^{\prime}\in E(g_{n})\setminus\{e\}$ is an element of ${\mathbb{F}}_{2}$ with $|g^{\prime}|_{\{a,b\}}\geq|g_{n}|_{\{a,b\}}$.
Since all of the $g_{n}$ are conjugate, they all have equal stable length. Therefore we can apply Theorem A.4 to get a degree 6 polynomial $P$ such that whenever $h_{1},\ldots,h_{m}\in B_{S}({|g_{n}|_{S}^{1/2}})$ and $|k_{1}|,\ldots,|k_{m}|\geq P(|g_{n}|_{S})$ we have $h_{1}g_{n}^{k_{1}}\cdots h_{m}g_{n}^{k_{m}}\neq e$.
We define $\phi_{N}:G*\langle z\rangle\to G$ by $\phi_{N}|_{G}=\operatorname{id}$ and $\phi_{N}(z)=g_{2CN^{2}}^{P(D(4CN^{2}+1))}$. We can now conclude as in the proof of Proposition 3.2. Clearly $\phi_{N}$ satisfies the subexponential growth condition, since $|g_{2CN^{2}}|_{S}\leq D|g_{2CN^{2}}|_{a,b}=D(4CN^{2}+1).$ It remains to show that $\phi_{n}$ is injective on $B_{S\cup\{z\}}(N)$. It suffices to show that
$h_{1}\ell_{1}\cdots h_{m}\ell_{m}\neq e$
for all $h_{i}\in G$ such that $|h_{i}|_{S}\leq 2N$ for all $1\leq i\leq m$ and $h_{i}\neq e$ for all $1<i<m$, and for all $\ell_{i}=\phi_{N}(z)^{p_{i}}$ where the $p_{i}$ are nonzero integers. But this follows from the previous paragraph’s assertion, since, taking $n=2CN^{2},$ we have $h_{i}\in B_{S}(2N)\subset B_{S}(\sqrt{(2n+1)/C})\subset B_{S}(|g_{n}|_{S}^{1/2})$ while the powers of $g_{n}$ are larger, in magnitude, than $P(D(2n+1))$ and thus are also larger than $P(|g_{n}|_{S})$. (Here, we assume $P$ is increasing, which is true for sufficiently large values.)